# On Space-Filling Input Design for Nonlinear Dynamic Model Learning: A Gaussian Process Approach

### v1

@@ Abstract
While optimal input design for linear systems has been well-established, no systematic approach exists for nonlinear systems, where robustness to extrapolation/interpolation errors is prioritized over minimizing estimated parameter variance. To address this issue, we develop a novel identification-oriented space-filling input design strategy for efficient covering of a well-defined region of interest. By placing a Gaussian Process (GP) prior on the joint input-state space, the proposed strategy leverages the GP posterior variance to construct a space-filling promoting input design cost function from the Bayesian point of view. This accordingly enables maximization of the coverage in the region of interest, thereby facilitating the generation of informative datasets. Furthermore, we theoretically prove that minimization of the cost function indicates the space-filling behavior. The effectiveness of the proposed strategy is demonstrated on both an academic example and a mass-spring-damper system, highlighting its potential practical impact on efficient exploration in identification scenarios for nonlinear dynamical systems.
@@ I Introduction
Nonlinear system identification [1] is a well-established topic, and has been applied in a broad range of domains such as mechanics and physics [2, 3]. In general, the quality of the model derived from identification is substantially dependent on the data collected during the identification procedure and input design plays a significant role in ensuring data informativity. In the nonlinear case, this especially concerns the extent to which the measurement dataset achieves effective data coverage of the input, output and state space in the region of interest, ensuring a comprehensive exploration of the dynamical system. However, such an input design for dynamical systems is challenging because the current output depends on the past system inputs and outputs.
@@
The overall process aiming at overcoming these challenges is often called identification-oriented input design and its objective is that given an expected modeling accuracy, obtain the information required to model the system dynamics at the lowest experimental cost [1]. There has been extensive research conducted on identification-oriented input design for linear systems [4, 5, 6]. State-of-the-art methods minimize the variance on the parameter estimate of an assumed parametric model of the system by maximizing the Fisher information matrix $\mathcal{M}$. Such input design has been considered both from the nominal and robust point of view to optimize some forms of $\mathcal{M}$ (D-optimality [7], A-optimality [8], E-optimality [9]). In the linear system case, this problem effectively translates into a convex optimization, where the solution is given by an optimal design of the power spectrum of the input signal [10].
@@
In contrast to the input design for linear systems, there is no general systematic framework available to address the input design problem for a wide range of nonlinear systems due to their dependence on both frequency-domain characteristics and time-domain evolution. Solution methods exist only for simple nonlinear systems, such as Hammerstein and Wiener [11]. Furthermore, given that the dominant source of errors in nonlinear system identification is most often the generalization error, i.e., model errors on regions of the operating space that are not contained in the estimation dataset [1], input design that minimizes the variance of the parameter estimate is not justified. Hence, an input design approach should be adopted to minimize the model error over a considered range of operations of the system of interest.
@@
A space-filling input ensures that the model errors are well-behaved and that the estimated model using this dataset would generalize well over the full region of interest. Various methods [12, 13, 14, 15] have been explored for space-filling input design in nonlinear systems, and it has also been shown that accurate model estimation can be obtained with the space-filling dataset in a set-membership setting [16]. Existing space-filling promoting cost functions and algorithms [17, 13, 18] mainly penalize the mutual distance between data points. While substantial work has been done on exploring various space-filling promoting cost functions, they are typically limited to some specific input classes (e.g. piecewise constant inputs [12, 14, 13]), and do not consider constraints on the input (e.g. spectral constraints). Additionally, active learning, which shares a similar spirit with input design, is used to actively explore the space by selecting sampling points that maximize the variance-based cost [19] or information gain [20]. Many active learning approaches follow a MPC-like procedure where the next sampling point (input sequence in MPC) is actively selected based on a specific information criterion. In contrast, space-filling input design, as an experiment design method, emphasizes the thorough coverage of the entire space by solving an offline design problem, without the need for state measurements of the system. Similarly to MPC-based methods, a model of the system needs to be available.
@@
In this letter, we study how to efficiently generate an informative dataset that can maximize the space coverage for a fixed length of the experiment. To this end, we propose a simple, yet effective identification-oriented Gaussian process (GP)-based space-filling input design strategy, where a GP prior is placed on the joint input-state space. The underlying philosophy of the proposed method is to solve the filling problem by minimizing the total posterior variance w.r.t. the user-selected anchor points distributed in a region of interest. The main contributions of this letter are summarized as:
@@
The remainder of this letter is organized as follows. We describe the problem formulation in Section II and summarize the background in Section III. The proposed GP-based space-filling input design method is detailed in Section IV. Experiment results are given in Section V, followed by the conclusions in Section VI.
@@ II-A Data-generating system
where $u(k)\in\mathcal{U}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$, $x(k)\in\mathcal{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$, and $y(k)\in\mathcal{Y}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$ are the input, state and output signals of the system at time instant $k\in\mathbb{Z}$, respectively, and $\mathcal{Z}:\mathcal{U}\times\mathcal{X}$ denotes the joint input-state space. The smooth nonlinear functions $f(\cdot):\mathcal{Z}\to\mathcal{X}$ and $g(\cdot):\mathcal{Z}\to\mathcal{Y}$ represent the known model of the system with $f\in\mathcal{C}_{1}$.
@@
The state-space representation (1) is controllable, which indicates that for any two states $x_{1},x_{2}\in\mathcal{X}$ there exists a finite time $T_{\mathrm{d}}\in\mathbb{Z}$ and an admissible control function $u:[0,T_{\mathrm{d}}]\to\mathcal{U}$ such that $x_{2}=\phi(T_{\mathrm{d}},0,x_{1},u)$, where $\phi(t,0,x_{0},u)$ denotes the unique solution of (1) at time $t\geq 0$ for a particular input function $u(\cdot)$ and initial condition $x(0)=x_{0}$.
@@ II-B Space-Filling Input Design Problem
It is worth noting that the optimal input design problem corresponds to a chicken-egg dilemma because to optimally design an input signal for which the output response will be informative w.r.t. the identification objective, perfect knowledge of the underlying system dynamics in (1) is necessary, while in order to completely capture the dynamics of (1) from data, we need a data set that is informative w.r.t. our model structure. This inherent interdependency leads to a challenge in effectively addressing both tasks concurrently within the system identification cycle. In this letter, we mainly consider the space-filling input design problem under a known nonlinear system model (1) irrespectively that it is the full or only the currently available model knowledge about the system to be identified.
@@
Under the condition that the model is unknown or it is approximately known with uncertainties, an iterative procedure can be adopted. Specifically, one may begin by assuming a region to explore and perform identification based on the dataset collected within this region. Subsequently, using the refined model, the space-filling input design can be repeated, taking into account the previously measured datasets. This alternating procedure continues until it achieves an adequate level of modeling accuracy.
@@
$\mathscr{U}(\theta):={\{\theta\in{\Theta},k\in\mathbb{T}|u(k;\theta):\mathbb{T}\times\Theta\to\mathbb{R}^{n_{\mathrm{u}}}\}}$ (2)
where $\mathbb{T}$ is a finite interval in $\mathbb{Z}$ and satisfies the condition:
@@
    $\mathscr{U}_{\text{ms}}(\theta)\!:=\!\left\{\!{k\!\in\!\mathbb{T}}\bigg{|}u(k;\theta)\!=\!\sum_{l=0}^{N\!_{f}\!-\!1}\!A_{l}\!\sin\!\left(\!2\pi l\frac{f_{0}}{f_{s}}\!k\!+\!\phi_{l}\!\right)\!\right\}\!$ (3)
@@
    $\mathscr{U}_{\text{pw}}(\theta)\!:=\!\left\{\!{k\!\in\!\mathbb{T}}\bigg{|}u(k;\theta)\!=\!\sum_{l=0}^{N\!_{p}-1}\!A_{l}\delta\!\left(p_{l-1},p_{l},k\right)\!\right\}$ (4)
@@
    $\delta(p_{l-1},p_{l},k)=\begin{cases}1,&p_{l-1}\leq k<p_{l}\\ 0,&\text{otherwise}\end{cases}$ (5)
    denotes the duration of each amplitude level with dwell time $p_{l}\in\mathbb{T}$.
As discussed in Section I, a space-filling input facilitates an identified model with a small error, allowing the model to represent well over the region of interest, which is defined as:
@@
The primary goal of this letter is to design the parameter vector $\theta\in{\Theta}$ such that when the input signal $u(k;\theta)$ is applied to the assumed, known system (1), the resulting dataset is expected to achieve an effective space-filling property, ensuring a good coverage within $\tilde{\mathcal{Z}}$.
@@ III Gaussian Process Models
In this section, we introduce some basic concepts from GP modeling. GP regression [21] is a well-known tool for constructing nonparametric, probabilistic models directly from data, and allows for flexible specification of prior assumptions on the system. To construct a hypothetical model of $f:\mathcal{Z}\to\mathcal{X}$, we consider $h:\mathcal{Z}\to\mathcal{X}$ as a GP. In terms of definition, a Gaussian Process $\mathcal{GP}:\mathbb{R}^{n_{\mathrm{z}}}\to\mathbb{R}$ assigns to every point ${z}\in\mathbb{R}^{n_{\mathrm{z}}}$ a random variable ${h}({z})$ taking values in $\mathbb{R}$ such that, for any finite set of points $\{{z}_{\tau}\}_{\tau=1}^{N}\subset\mathbb{R}^{n_{\mathrm{z}}}$, the joint probability distribution of $\mathcal{GP}\left({z}_{1}\right),\ldots,\mathcal{GP}\left({z}_{N}\right)$ is Gaussian. A GP can be fully characterized by its mean function $\mu(z):\mathcal{Z}\to\mathbb{R}$ and its covariance function $\kappa(z,z^{\prime}):\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ and it is denoted as: $h(z)\sim\mathcal{GP}\left({\mu}({z}),{\kappa}\left({z},{z}^{\prime}\right)\right)$.
Given a training dataset $\mathcal{D}=\{Z,H\}$ with the training input matrix $Z=\mathrm{vec}(z_{1},...,z_{N})\in\mathbb{R}^{N\times n_{\mathrm{z}}}$ and the training output $H=\mathrm{vec}(h_{1},...,h_{N})\in\mathbb{R}^{N\times 1}$, with $h_{i}=h(z_{i})$. Under the assumption that function $h(z)$ follows a GP, one can define the Gaussian prior distribution for $h$ as $\mathbb{P}\left(H\right)=\mathcal{N}(H\mid 0,K_{N})$ where $K_{N}\in\mathbb{R}^{N\times N}$ is the Gram matrix with entries $[K_{N}]_{i,j}=\kappa(z_{i},z_{j})$. Given a query point $\bar{z}\in\mathbb{R}^{n_{\mathrm{z}}}$, the joint distribution of $H$ and $\bar{h}(\bar{z})$ is:
@@
Then conditioning the joint distribution on $\mathcal{D}$ leads to the predictive distribution of $\bar{h}(\bar{z})$ given an input $\bar{z}$:
$\mathbb{P}(\bar{h}\mid\bar{z},\mathcal{D})=\mathcal{N}(\bar{h}\mid\mu(\bar{h}\mid\bar{z},\mathcal{D}),\mathrm{Var}(\bar{h}\mid\bar{z},\mathcal{D}))$ (7)
with the mean and variance expressed respectively as:
@@
According to (8b), it can be deduced that the posterior variance evaluated at $\bar{z}$ only depends on the GP inputs $Z$, which corresponds to the particular property of Gaussian distribution. Hence, the posterior variance of a GP inherently represents the distribution of GP inputs $Z$ in the given dataset over the space $\mathcal{Z}$.
@@ IV GP-based Space-filling Input Design
In this section, we propose a GP-based space-filling input design algorithm, enabling the acquisition of an informative dataset for data-driven model learning that achieves good coverage within the region of interest $\tilde{\mathcal{Z}}$. The main philosophy behind this is to firstly place a GP prior on a latent hypothetical model $\hat{x}(k+1)=\hat{f}(x(k),u(k))$ of the data-generating system (1), and then construct the space-filling promoting cost function from the Bayesian point of view.
@@ IV-A Classical Space-filling Design
A function ${d}(x_{1},x_{2}):\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{0}^{+}$ is a distance-based metric if it satisfies:
  • $d(x_{1},x_{2})=d(x_{2},x_{1})$, $\forall{x_{1},x_{2}}\in\mathcal{X}$,
  • $d(x_{1},x_{2})\geq 0$ with equality if and only if $x_{1}=x_{2}$,
  • $d(x_{1},x_{2})\leq d(x_{1},x_{3})+d(x_{2},x_{3})$, $\forall{x_{1},x_{2}}$ and $x_{3}\in\mathcal{X}$.
Definition 3
Consider a dataset $\mathcal{D}_{N}=\{z_{i}\}_{i=1}^{N}$ and a region of interest ${\tilde{\mathcal{Z}}}$. Let $d(\varsigma,\mathcal{D}_{N})=\min_{z_{i}\in\mathcal{D}_{N}}d(\varsigma,z_{i})$ denote the minimum distance w.r.t. $\varsigma$ where $d(\cdot,\cdot)$ is given in Definition 2. Then
@@
is the so-called filling distance indicating the radius of the largest ball that can be placed in $\tilde{\mathcal{Z}}$ which does not contain any point in $\mathcal{D}_{N}$. Specifically, if there exists a constant $\epsilon$ such that ${\min_{\mathcal{D}_{N}}\rho({\mathcal{D}_{N}})}<\epsilon,~{}\epsilon>0$, then $\mathcal{D}_{N}$ is said to have the at least $1/\epsilon$ density w.r.t $\tilde{\mathcal{Z}}$, i.e., the $\epsilon$-space filling property.
The classical space-filling design often focuses on the distribution of the points within $\mathcal{D}_{N}$, without considering the temporal dynamics between the points, that is, the underlying dynamics that generate these data points are neglected. Hence, we highlight the nonlinear dynamical systems and design identification-oriented space-filling input signals.

### v2

@@ Abstract
While optimal input design for linear systems has been well-established, no systematic approach exists for nonlinear systems where robustness to extrapolation/interpolation errors is prioritized over minimizing estimated parameter variance. To address this issue, we develop a novel space-filling input design strategy for nonlinear system identification that ensures data coverage of a given region of interest. By placing a Gaussian Process (GP) prior on the joint input-state space, the proposed strategy leverages the GP posterior variance to construct a cost function that promotes space-filling input design. Consequently, this enables maximization of the coverage in the region of interest, thereby facilitating the generation of informative datasets. Furthermore, we theoretically prove that minimization of the cost function implies the space-filling property of the obtained data. Effectiveness of the proposed strategy is demonstrated on both an academic and a mass-spring-damper example, highlighting its potential practical impact on efficient exploration of the dynamics of nonlinear systems.
@@ I Introduction
When performing nonlinear system identification [1], the quality of the model strongly depends on the quality of the data. In general, experiment design aims to obtain the information required to model system dynamics at the lowest experimental cost, while achieving expected modeling accuracy and respecting possible constraints to the experiment itself [1, 2]. In-depth research has been conducted on input design for linear time-invariant (LTI) systems [3, 4, 5]. State-of-the-art methods minimize the variance on the parameter estimate of an assumed parametric model of the system by maximizing the Fisher information matrix $\mathcal{M}$, under nominal and robust settings by D-, A, or E-optimality criterion. In the LTI case, this leads to convex optimization problems, giving optimal design of the power spectrum of the input signal [6]. Other approaches focus explicitly on the intended application of the model, known as application-oriented input design (see [2, 7, 4], etc.), which often links the design criteria to a specific use case such as control performance.
In contrast to the linear case, nonlinear systems generally require the input design to account for both spectral and time-domain characteristics, such as amplitude distribution and temporal evolution due to their inherently coupled behavior. Most existing methods target specific classes of nonlinear systems, such as Hammerstein and Wiener such as [8]. Others consider a wider class of systems but are limited to an input sequence with a finite number of different values [9], or aim to minimize the mutual coherence of the regressor in the sparse nonlinear system identification problem [10].
@@
Given that the dominant source of errors in nonlinear system identification is most often the generalization error, i.e., model error over a region of the operating space that is not contained in the estimation dataset [1], input design that minimizes the variance of the parameter estimate is not justified. Instead, the design should focus on minimizing the model error over an operating range of interest.
@@
A space-filling input ensures that, after identification, the model errors are well behaved and that the estimated model generalizes well over the full region of interest. Various methods [11, 12, 13, 14, 15, 16] have been explored for the design of space-filling inputs for nonlinear systems, utilizing, for instance, distance-based criteria that penalize mutual proximity between data points [11, 12, 13], as well as more recent approaches based on Jensen-Shannon divergence [15] or frequency-aware space-filling costs [16]. However, most existing approaches are restricted to specific input classes (e.g., piecewise constant inputs [17, 18, 12]) and do not consider constraints on the input (e.g., spectral constraints). In addition, some recent methods require matching the empirical data distribution to a fixed target distribution, which may be unnecessarily restrictive, or rely on non-smooth cost functions that limit the use of gradient-based optimization. Active learning, which aims at similar objectives as input design, can also be used to actively explore the space of interest by iteratively selecting sampling points that, based on the current model estimate, maximize a variance-based cost [19] or the information gain [20]. In contrast, space-filling input design, as an experiment design method, focuses on thorough coverage of the entire space by solving an offline design problem using a prior model of the system without requiring real-time state information during data collection.
@@
In this paper, we study how to efficiently generate an informative dataset that can maximize the space coverage for a fixed length of the experiment. To this end, we propose a simple, yet effective Gaussian process (GP)-based space-filling input design strategy. The underlying philosophy of the proposed method is to solve the filling problem by minimizing the total posterior variance w.r.t. some user-selected anchor points distributed in a region of interest. The main contributions are summarized as:
@@
The paper is organized as follows. The problem formulation is described in Section II and GPs are introduced in Section III. The proposed GP-based space-filling input design method is detailed in Section IV. Examples are given in Section V, followed by the conclusions in Section VI.
@@ II-A Data-generating system
where $u(k)\in\mathcal{U}\subseteq\mathbb{R}^{n_{\mathrm{u}}}$, $x(k)\in\mathcal{X}\subseteq\mathbb{R}^{n_{\mathrm{x}}}$, and $y(k)\in\mathcal{Y}\subseteq\mathbb{R}^{n_{\mathrm{y}}}$ are the input, state and output signals of the system at time instant $k\in\mathbb{Z}$, respectively, and $\mathcal{Z}:\mathcal{U}\times\mathcal{X}$ denotes the joint input-state space. The smooth nonlinear functions $f(\cdot):\mathcal{Z}\to\mathcal{X}$ and $g(\cdot):\mathcal{Z}\to\mathcal{Y}$ represent the known model of the system, with $f$ assumed to be continuously differentiable.
@@
[21] The state-space representation (1) is controllable, which indicates that for any two states $x_{1},x_{2}\in\mathcal{X}$ there exists a $T_{\mathrm{d}}\in\mathbb{Z}_{>0}$ and an input sequence $u:\{0,1,...,T_{\mathrm{d}}-1\}\to\mathcal{U}$ such that $x_{2}\!=\!\phi(T_{\mathrm{d}},0,x_{1},u)$, where $\phi(k,0,x_{0},u)$ denotes the unique solution of (1) at time step $k\geq 0$ for a particular $u$ and initial condition $x(0)\!=\!x_{0}$.
@@ II-B Space-Filling Input Design Problem
It is worth noting that the optimal input design problem corresponds to a chicken-egg dilemma: to optimally design an input signal that results in the informative output response w.r.t. the identification objective, perfect knowledge of the underlying system dynamics in (1) is necessary; while to completely capture the dynamics of (1), we need a data set that is informative w.r.t. our model structure. This inherent interdependency makes it challenging to effectively address both tasks concurrently within the identification cycle. In this paper, we focus on the space-filling input design problem under a known nonlinear system model (1) regardless of whether it represents full or partial knowledge of the system.
@@
If the model is unknown, or it is approximately known with uncertainties, an iterative procedure can be adopted: one can design an experiment on the uncertain model, or start with a generic experiment, and estimate a model from this initial dataset. Subsequently, using the refined model, the space-filling input design can be repeated, taking into account the previously measured datasets until it achieves an adequate level of modeling accuracy.
@@
$\mathscr{U}(\Theta):={\{u(k,\theta):\mathbb{T}\times\Theta\to\mathbb{R}^{n_{\mathrm{u}}}\}}$ (2)
where $\mathbb{T}$ is a finite interval in $\mathbb{Z}$ and (2) satisfies the condition:
@@
    $\mathscr{U}_{\text{ms}}(\Theta)\!:=\!\left\{\!u(k,\theta)\!=\!\!\sum_{l=0}^{N\!_{f}\!-\!1}\!A_{l}\!\sin\!\left(\!2\pi l\frac{f_{0}}{f_{s}}\!k\!+\!\phi_{l}\!\right)\!\bigg{|}\theta\!\in\!\Theta\!\right\}\!$ (3)
@@
    $\mathscr{U}_{\text{pw}}(\Theta)\!:=\!\left\{\!u(k,\theta)\!=\!\!\!\sum_{l=0}^{N\!_{p}-1}\!A_{l}\delta\!\left(p_{l},p_{l+1},k\right)\!\bigg{|}\theta\!\in\!\Theta\!\right\}$ (4)
@@
    $\delta(p_{l},p_{l+1},k)=\begin{cases}1,&p_{l}\leq k<p_{l+1}\\ 0,&\text{otherwise}\end{cases}$ (5)
    denotes the duration of each amplitude level with switching instances $p_{l}\in\mathbb{T}$.
As discussed in Section I, a space-filling input allows the model to represent the system well over the region of interest, which is defined as:
@@
The primary goal of this paper is to design the parameter vector $\theta\in{\Theta}$ such that when the input signal $u(k,\theta)$ is applied to the assumed, known system (1), the resulting dataset is expected to achieve an effective space-filling property, ensuring a good coverage of $\tilde{\mathcal{Z}}$.
@@ III Gaussian Process Models
In this section, we introduce some basic concepts of GP modeling. GP regression [22] is a well-known tool for constructing nonparametric, probabilistic models directly from data, and allows for flexible specification of prior assumptions on the system. To construct a hypothetical model of $f:\mathcal{Z}\to\mathcal{X}$, we consider $h:\mathcal{Z}\to\mathcal{X}$ as a GP. In terms of definition, a Gaussian Process $\mathcal{GP}:\mathbb{R}^{n_{\mathrm{z}}}\to\mathbb{R}$ assigns to every point ${z}\in\mathbb{R}^{n_{\mathrm{z}}}$ a random variable ${h}({z})$ taking values in $\mathbb{R}$ such that, for any finite set of points $\{{z}_{\tau}\}_{\tau\!=\!1}^{N}\subset\mathbb{R}^{n_{\mathrm{z}}}$, the joint probability distribution of $h\left({z}_{1}\right),\ldots,h\left({z}_{N}\right)$ is Gaussian. A GP can be fully characterized by its mean function $\mu(z):\mathcal{Z}\to\mathbb{R}$ and its covariance function $\kappa(z,z^{\prime}):\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ and it is denoted as: $h(z)\sim\mathcal{GP}\left({\mu}({z}),{\kappa}\left({z},{z}^{\prime}\right)\right)$.
Given a training dataset $\mathcal{D}\!=\!\{Z,H\}$ with the training input matrix $Z\!=\!\mathrm{vec}(z_{1},...,z_{N})\in\mathbb{R}^{N\times n_{\mathrm{z}}}$ and the training output $H\!=\!\mathrm{vec}(h_{1},...,h_{N})\in\mathbb{R}^{N\times 1}$, with $h_{i}\!=\!h(z_{i})$. Under the assumption that the function $h(z)$ follows a GP, one can define the Gaussian prior distribution for $h$ as $\mathbb{P}\left(H\right)\!=\!\mathcal{N}(H\mid 0,K_{N})$ where $K_{N}\in\mathbb{R}^{N\times N}$ is the Gram matrix with entries $[K_{N}]_{i,j}\!=\!\kappa(z_{i},z_{j})$. Given a query point $\bar{z}\in\mathbb{R}^{n_{\mathrm{z}}}$, let $\bar{h}\!:=\!h(\bar{z})$ denote the random output at $\bar{z}$. Then, the joint distribution of $H$ and $\bar{h}$ is:
@@
Then conditioning the joint distribution on $\mathcal{D}$ leads to the predictive distribution of $\bar{h}$ given an input $\bar{z}$: $\mathbb{P}(\bar{h}\mid\bar{z},\mathcal{D})\!=\!\mathcal{N}(\bar{h}\mid\mu(\bar{h}\mid\bar{z},\mathcal{D}),\mathrm{Var}(\bar{h}\mid\bar{z},\mathcal{D}))$ with the mean and variance expressed respectively as:
@@
According to (7b), the posterior variance evaluated at $\bar{z}$ only depends on the GP inputs $Z$. Hence, the posterior variance of a GP inherently represents the distribution of GP inputs $Z$ in the given dataset over $\mathcal{Z}$.
@@ IV GP-based Space-filling Input Design
In this section, we propose a GP-based space-filling input design algorithm, enabling the acquisition of an informative dataset for data-driven model learning that achieves good coverage within the region of interest $\tilde{\mathcal{Z}}$. The main philosophy behind this is to firstly place a GP prior on a latent hypothetical model $\hat{x}(k+1)\!=\!\hat{f}(x(k),u(k))$ of the data-generating system (1), and then construct a space-filling promoting cost function from the Bayesian point of view.
@@ IV-A Classical Space-filling Design
Consider a dataset $\mathcal{D}_{N}\!=\!\{z_{i}\}_{i\!=\!1}^{N}$ and a region of interest ${\tilde{\mathcal{Z}}}$. Let $d(\varsigma,\mathcal{D}_{N})\!=\!\min_{z_{i}\in\mathcal{D}_{N}}d(\varsigma,z_{i})$ denote the minimum distance w.r.t. $\varsigma$ where $d(\cdot,\cdot):\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}_{0}^{+}$ is a user-chosen metric (e.g. Euclidean distance). Then
@@
is the so-called filling distance indicating the radius of the largest ball that can be placed in $\tilde{\mathcal{Z}}$ which does not contain any point in $\mathcal{D}_{N}$. Specifically, if there exists a constant $\epsilon$ such that ${\rho({\mathcal{D}_{N}})}<\epsilon,~{}\epsilon>0$, then $\mathcal{D}_{N}$ is said to have the at least $1/\epsilon$ density w.r.t $\tilde{\mathcal{Z}}$, i.e., the $\epsilon$-space filling property.
In contrast to this work, classical space-filling design only focuses on the distribution of the points within $\mathcal{D}_{N}$, without considering the temporal dynamics between the points, i.e., the underlying dynamics that generate these data points.